OpenStudy (anonymous):

How do I convert r = 3cot theta csc theta to Cartesian Form

OpenStudy (anonymous):

\[\cot(\theta)=1/tg(\theta)=1/(y/x)=x/y\]\[\csc(\theta)=1/\sin(\theta)=1/(y/r)=r/y\]Then you have that:\[r=3\frac{ x }{ y }·\frac{ r }{ y } \rightarrow \frac{ 3x }{ y^2}=1\rightarrow y=\sqrt{3x}\]

OpenStudy (anonymous):

The correct way to express the solution is: \[3x-y^2=0\]

OpenStudy (anonymous):

Thank you so much!! This was really helpful.

OpenStudy (anonymous):

you are welcome